OpenStudy (anonymous):

The difference of two numbers is 1 What is the smallest possible value for the sum of their squares?

OpenStudy (google):

x-y=1 Minimize x^2+y^2. do the same thing as before, just be careful

OpenStudy (anonymous):

The two numbers are 1/2 and -1/2.

OpenStudy (vishal_kothari):

make it symmetrical with respect to 0. 1/2 - (-1/2) = 1 (-1/2)^2 + (1/2)^2 = 1/2...

OpenStudy (anonymous):

Vishal where you from?